OpenStudy (anonymous):

The length of a rectangle is 4 centimeters less than twice its width. The perimeter of the rectangle is 34 cm. What are the dimensions of the rectangle?

OpenStudy (kewlgeek555):

Width = w Length = w-4 Perimeter = 34 34 = ((w-4) * 2) + (w * 2) So based on my calculations w = 10.5 So the length is 10.5 - 4 = 6.5 Let's see if this is correct. 34 = ((10.5 - 4) * 2) + 10.5 * 2) 34 = (6.5 * 2) + (10.5 * 2) 34 = 13 + (10.5 * 2) 34 = 13 + 21 34 = 34

OpenStudy (kewlgeek555):

Also: Welcome to OpenStudy. c:

OpenStudy (anonymous):

thanks:)

OpenStudy (kewlgeek555):

You are welcome. Also, when someone answers a question, you can give them a "medal" by pressing the blue "best response" button. ;D

OpenStudy (kewlgeek555):

(if you need anymore help with the website, just message me. c;) You can do that by hovering over my profile picture and pressing "send message" You can also make friends in the chats below - they should be either orange or green. ;D Good luck and have a nice day. :3